Show that the square of an even number is an even number using a direct proof.



1
Expert's answer
2022-06-13T14:30:50-0400

Given "n" is even. Then "\exist k\in \Z" such that "n=2k."

Then


"n^2=(2k)^2=4k^2=2(2k^2)"

Let "m=2k^2, k\in \Z." Then "m=2k^2, m\in \Z."

So "n^2=2m, m\in \Z." This means that "n^2" is even.

Therefore the square of an even number is an even number.